Alyeski Tours operates day tours of coastal glaciers in Alaska on its tour boat the Blue Glacier. Management has identified two cost drivers—the number of cruises and the number of passengers—that it uses in its budgeting and performance reports. The company publishes a schedule of day cruises that it may supplement with special sailings if there is sufficient demand. Up to 84 passengers can be accommodated on the tour boat. Data concerning the company’s cost formulas appear below

Fixed Cost per Month Cost per Cruise Cost per Passenger
Vessel operating costs $ 6,700 $ 476.00 $ 3.10
Advertising $ 3,000
Administrative costs $ 5,900 $ 38.00 $ 1.50
Insurance $ 3,700

For example, vessel operating costs should be $6,700 per month plus $476.00 per cruise plus $3.10 per passenger. The company’s sales should average $29.00 per passenger. In July, the company provided 53 cruises for a total of 3,000 passengers.Required:

Prepare the company’s flexible budget for July.

Alyeski Tours
Flexible Budget
For the Month Ended July 31
Revenue
Expenses:
Vessel operating costs
Advertising
Administrative costs
Insurance
Total expense
Net operating income
Answer #1
Alyeski Tours
Flexible Budget
For the Month Ended July 31
Actual cruises 53
Actual passengers 3000
Revenue 87000 =3000*29
Expenses:
Vessel operating costs 41228 =6700+(476*53)+(3000*3.10)
Advertising 3000
Administrative costs 12414 =5900+(38*53)+(3000*1.50)
Insurance 3700
Total expense 60342
Net operating income 26658